Hayhayz (hayhayz):

What are the coordinates of the turning point for the function f(x) = (x + 2)3 − 4? @bibby

OpenStudy (trojanpoem):

Derive the function twice. Is it f(x) = (x+2)^3 - 4 ?

Hayhayz (hayhayz):

Yes

OpenStudy (trojanpoem):

Do you know how to find the derivative for a function ?

Hayhayz (hayhayz):

No...

OpenStudy (trojanpoem):

To find the derivative of an exponential function with constant power: y = const => y' = 0 y = (x)^n => y' = n (x)^(n-1) * x' y = cosu => y' = -u' sinu y = sinu => y' = u' cosu There is a lot more, inverse function , logarithms , variable power, etc. (critical points) maxima , minima : first derivative. turning point, concavity : second derivative So let's back to the question: f(x) = (x+2)^3 - 4 f'(x) = 3(x+2)^2 - 0 ( -4 is a constant so its derivative is zero) f'(x) = 3(x+2)^2 f''(x) = 6 (x+2) = 6x + 12 Let f''(x) = 0 6x = -12 x = -2 ( turning point x coordinate) Plug it into the original function to find the y coordinate.

Hayhayz (hayhayz):

y would be -4?

OpenStudy (trojanpoem):

Exactly !

OpenStudy (trojanpoem):

You could also find it using graphing.

Hayhayz (hayhayz):

Thanks !!

OpenStudy (trojanpoem):

Any time.